Two students are on a balcony 19.1 m above the str
Answer #1

given initial velocity u = 13.8 m/s

h = 19.1 m

1) first ball

v^2 - u^2 = 2 * a * s

v^2 - 13.8^2 = 2*(-9.8) * (-19.1)

v = 23.76 m/s

2) for second ball

when it misses the window it has same initial velocity 13.8 m/s

so when it strikes the ground

v = 23.76 m/s

3)

time for first ball is from

s = ut + (1/2 * a*t^2)

-19.1 = -13.8 t - (0.5*(9.8)*(t)^2

t = 1.01 s

for second ball

s = ut + (1/2 * a*t^2)

-19.1 = 13.8 t - (0.5*(9.8)*(t)^2

t = 3.83 s

difference is 3.83 - 1.01 = 2.82 s

4)

for first ball

s = ut + (1/2 * a*t^2)

s1 = -(13.8 * 0.598) - (0.5*(9.8)*(0.598)^2
s1 = 10 m ( down ward)

for second ball

s2 = (13.8 * 0.598) - (0.5*(9.8)*(0.598)^2

s2 = 6.5 m ( upward)

they are apart by a distance of s = 10 + 6.5 m

s = 16.5 m